A merry-go-round full of kids initially has 5.33


#3
Answer #1

3 ans :

step1:

the time taken by rolling body to reach the height

t=1/sin(theta){2h/g(1+B)}^1/2

step2:

the B value for

solid ball =2/5

disk=1/2

hoop =1

step3:

now use the step1&step2 we get each one

time taken for solid ball to reach height =>Tball={2h/g(1+B)}^1/2

={2/9.8(1+2/5)}^1/2

={0.2041*1.4}^1/2

=0.535 sec

time taken for hoop to reach height =>Thoop={2/9.8(1+1)}^1/2

=0.64 sec

time taken for disk to reach height =>Tdisk={2/9.8(1+1/2)}^1/2

=0.6 sec

step4:

basing on step3

we decided the correct option is 3